Q12. a) Suppose the given confidence interval estimates the true population mean as 3.5<μ<13.1 with a 95% level of confidence when σ is known.

(i) Find the point estimates for the unknown population mean.

(ii) Find the margin of error.

(iii) Give the interpretation of the confidence interval.

(iv) Name two ways of decreasing the width of the confidence interval.

b) State the assumptions necessary for linear regression model Y=A+Bx+ε

c) Let p^ be a sample proportion based on a random sample of size n from a binomial distribution with unknown pp, such that γp^>5 and γq>5.

Construct a (1−α)100% confidence interval for the unknown population proportion

p using p^≈N

(μp^=pp^≈N(μp^=p, σp^=PVn−−−√)
Answer #1

Given, (1-4) 100 = 95 I-a -95 100 | - = 0.gs 2= 1-0.95 a=005 il The point estimates for the unknown population mean. is 3.5 <E = 2 ( 0 ) X + upper limit - 13. 1 = 8.30 + E E = 1351 - 836 E = 4.80 ... The margin of of error .is 4.80 (iii) We estimateE = 2 ( 0 ) X + upper limit - 13. 1 = 8.30 + E E = 1351 - 836 E = 4.80 ... The margin of of error .is 4.80 (iii) We estimate